OpenStudy (anonymous):

the area of an isosceles right triangle is 128 units^2 what is the perimeter of right trangle

OpenStudy (saifoo.khan):

is there anything more given?

OpenStudy (anonymous):

no thats it what do you need

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

 Area here is equal to (1/2)*base*height =(1/2)*x*x=(x^2)/2 But area is equal to 128 units^2. (x^2)/2=128 x=16 From the triangle h^2=x^2+x^2 h=sqrt(2*x^2) h=xsqrt2 Perimeter=x+x+h =16+16+16sqrt2=32+16sqrt2